# JudgeAgent: Knowledge-wise and Dynamic LLM Evaluation with Agent-as-Interviewer

### v3

@@ Appendix B Analysis and Proofs of Adaptive Difficulty-Control Mechanisms
$\textbf{C1.}\quad avg_{s}^{\prime}=\frac{avg_{s}N+c_{it}}{N+1}>avg_{s},\quad\text{holds true always when}\ \ avg_{s}\leq t_{(i,i+1)}$ (2)
$\textbf{C2.}\quad\exists\ avg_{s0},N_{0},\quad\text{s.t.}\ avg_{s}^{\prime}=\frac{avg_{s0}N_{0}+c_{it}}{N_{0}+1}>t_{(i,i+1)}$
@@ Appendix C Construction of Context Graph
$N=(e\,,\,\mathcal{C}\,,\,\mathcal{D})\ ,\ \text{where}\ \mathcal{C}=\{c_{1},c_{2}...\}\ ,\ \mathcal{D}=\{d_{1},d_{2},$...\} (6)
@@ D.1 Details of Datasets
MultiHop-RAG(Tang & Yang, 2024) consists of phrase Q&A queries, their ground truth answers, and the associated supported evidence constructed from news articles published between September and December 2023. The background knowledge texts of MultiHop-RAG are provided as supporting evidence along with the questions.
D.2 Preprocessing of Datasets
To verify Agent-as-Interviewer’s ability to evaluate knowledge deficiencies in target models, we remove the background knowledge of the questions from MedQA and MultiHopRAG during evaluation. Otherwise, it would only test the target’s reading comprehension ability rather than knowledge deficiencies. In contrast, we provided the background text when using QuALITY for evaluation, as the questions in QuALITY are highly dependent on the text content, and many of the texts are fictional narratives. Therefore, we use QuALITY to validate the JudgeAgent’s effectiveness in guiding comprehension and reasoning rather than discovering knowledge deficiencies.
@@ Appendix E Additional Experiment Analysis
E.1 The Resilience Against Data Contamination
Can Agent-as-Interviewer mitigate the challenges of data contamination in static benchmarking paradigms? In this section, we simulate a scenario where the static benchmarking evaluation paradigm suffers from data contamination by deliberately exposing the evaluation questions to LLMs during their training process. We selected Llama3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct as base models, and constructed supervised fine-tuning (SFT) data from the MedQA and MultiHop-RAG benchmarks, which are intended for evaluation in the main experiments. These training data were used to fine-tune the selected base models. By comparing the performance differences between the original base models and the fine-tuned models on both the static benchmark questions and the extended questions generated by JudgeAgent, we analyze and verify the resilience of the Agent-as-Interviewer paradigm and its derivative JudgeAgent against data contamination.
Table 5: The results for validating the resilience against data contamination. base means the performance on base static benchmark questions, and @K means extended questions at the K-th interation.
@@ E.1 The Resilience Against Data Contamination
In contrast, when evaluated on extended questions generated by JudgeAgent, the fine-tuned models and base models show little difference in performance. Additionally, fine-tuning even resulted in a decline on the MedQA benchmark. These results suggest that under the Agent-as-Interviewer dynamic evaluation paradigm, the generated questions maintain the validity of evaluation, even when the original questions have been exposed to the target model. Furthermore, under the same setting of LLM and benchmark, there is a marked gap between the performance gain on base questions and extended questions after fine-tuning, demonstrating that Agent-as-Interviewer exhibits considerable resilience to data contamination.
E.2 Cross Validation of the Evaluation Suggestions
Do the suggestions provided by Agent-as-Interviewer only take effect for the seed questions? To address this question, we designed cross-validation experiments from two perspectives.
First, considering that the evaluation suggestions are derived from a comprehensive evaluation of the responses to both the seed questions and their extended questions, we evaluated and compared the accuracy improvement of the suggestions on the extended questions versus the seed questions, aiming at verifying that the suggestions do not contain ”cheating information” specific to seed questions in the scenario after evaluation. The results are shown in Figure 5.
@@ E.2 Cross Validation of the Evaluation Suggestions
Secondly, we transferred the evaluation suggestions derived from seed questions to other questions with related knowledge concepts in the benchmark, to verify the effectiveness of the suggestions within the same knowledge domain. Specifically, we categorized questions based on the knowledge entities they contain, formalized as follows: given the context graph $\mathcal{G}$, for two questions $q_{1}$ and $q_{2}$ with knowledge entities $E_{1}=\{e|e\in q_{1}\land e\in\mathcal{G}\}$ and $E_{1}=\{e|e\in q_{2}\land e\in\mathcal{G}\}$, if $E_{1}$ and $E_{2}$ have a non-empty intersection, then $q_{1}$ and $q_{2}$ are considered related questions. In this experiment, the suggestions for a question were constructed from the suggestions of its related questions, excluding suggestions from the question itself, to assess and validate the transferability of the suggestions provided by the Agent-as-Interviewer. We screened out questions without relevant questions and those with only relevant questions based on non-knowledge entities, such as male, female, 2 years, etc. The results are shown in Table 6.
@@
First, we analyze the difference in the effectiveness of evaluation suggestions on seed questions versus extended questions. As shown in Figure 5, suggestions effectively improve performance for both seed questions and extended questions, with a relatively small gap in the degree of improvement. Notably, when Qwen3 and GPT-4.1 are used as target models, the first round of Qwen3 and the second round of GPT-4.1 exhibit even greater improvement than seed questions. These results indicate that although the suggestions only supplement knowledge for several key concepts, such as the case in Figure 7, such concise suggestions can still benefit both seed and extended questions, demonstrating that Agent-as-Interviewer is capable of identifying and addressing knowledge gaps in the target model, rather than simply providing ”cheating information” specific to seed questions.
@@
Next, we analyze the difference in the effectiveness of suggestions on questions that share the same knowledge concepts. As shown in Table 6, compared to their effectiveness on seed questions, the suggestions exhibit a slight decrease when applied to related questions, as indicated by a decline in CR and $\Delta$, and an increase in CtE. But the difference is minor, and the improvement remains notable, suggesting that the suggestions can be effectively transferred to other questions involving the same knowledge concepts. This further demonstrates that Agent-as-Interviewer can provide suggestions that do not simply serve as “cheating information” specific to seed questions.
However, the slight decline in evaluation effectiveness also indicates that the knowledge guidance provided in the suggestions is not fully aligned with the related questions. This may be because the overlapping entities between these related questions and seed questions do not correspond to core knowledge concepts. For example, suggestions centered on ”blood type” may be transferred to a question where ”serum” is the core knowledge concept, resulting in a partial mismatch.
The above experiment results demonstrate that the evaluation suggestions provided by Agent-as-Interviewer are not only applicable to seed questions but can also be transferred to other questions that share relevant core knowledge concepts.
E.3 Supplementary Parameter Analysis
What is the impact of batch size in the Benchmark Grading stage on the evaluations of the Agent-as-Interviewer? In Benchmark Grading stage, questions are divided into batches to comprehensively assess the target’s capabilities in a base level. These batches are also the basic units for question extension and evaluation feedback. Given a fixed rounds, the batch size is inversely proportional to the number of batches, extended questions, and evaluation suggestions, thereby influencing the time and resource consumption of the entire evaluation process. Can the batch size be maximized to reduce resource consumption while maintaining the effectiveness of evaluations? To address this question, we conducted a parameter analysis experiment, examining the evaluation effectiveness and time consumption under different batch sizes. The results are shown in Figure 6.
@@ Appendix G Prompts
We show the detailed prompt for Entity Extraction in Prompt 7, the prompt for generating extended questions in Prompt 8, the prompt for evaluating the performance of the target LLM in Prompt 9, and the prompt for querying the target LLM with suggestions in Prompt 10.
Prompt for Entity Extraction Please identify and label the entities in the following multiple sentences, and return the entity labeling results for each sentence. The results for each sentence should be independent, in JSON format, containing the sentence number, sentence text, and the list of recognized entities (including entity text, type, and position). Return format is a dictionary, with only one key ’labeled_data’, and the value is a list, each element is a dictionary containing the sentence text and the entity list. {{   ”labeled_data”:   [     {{”text”:”Sentence 1”, ”entity_list”: [{{”entity_text”: ””, ”entity_type”: ””}}]}},     {{”text”:”Sentence 2”, ”entity_list”: [{{”entity_text”: ””, ”entity_type”: ””}}]}},     …   ] }} Notice that ”text” should be only the sentence, not the whole article. Sentence list: { Sentences }
@@
Prompt for Generating Extended Questions As an interviewer, you are tasked with designing questions based on the provided texts. Your role involves crafting questions and correct answers that fulfill the following criteria: 1. **Focus on the Entity**: Ensure all questions consistently center around the specified entity from the article. 2. **Ensure Accuracy and Conciseness of Answers**: Verify that the provided answer is both correct for your designed question within the context and logic of the given text fragments, and ensure the answer is sufficiently concise—presented as a word or phrase, avoiding redundancy. 3. **Conform to difficulty requirements**: You need to design questions for the required difficulty levels, with specific requirements as follows:   (1). [easy]: **Encourage Knowledge Memorization**, design questions that assess whether respondents have memorized relevant knowledge. Create questions by directly extracting and blanking out content from the given passage.   (2). [medium]: **Encourage Knowledge Comprehension**: Design questions that prompt respondents to dissect and comprehend concepts involved in the topic. Avoid assessing only superficial knowledge retention.   (3). [hard]: **Encourage Knowledge Deep Analysis**: Design questions that prompt respondents to engage in deep thinking and analysis. Avoid merely testing knowledge recall or conceptual comprehension; do not simply extract fragments from the given passage to create fill-in-the-blank items. Encourage respondents to focus on entities within the question and employ logical skills for complex reasoning. Here are examples: { Examples } Now, given the following text fragments: { context } Based on the provided texts, please generate questions by following the requirements above and referencing the examples. Output in the specified JSON format below: {{   ”generated_question”:   [     {{       ”question”: ”Generated Question”,       ”answer”: ”Correct Answer of Generated Question”     }},   …   ] }}
@@
Prompt for Evaluation The following is the performance of an LLM in answering a series of questions: { list of questions, correct answers, and LLM’s answers } Please evaluate and analyze the interviewee’s performance based on the above performance using concise language from the following perspectives, and provide suggestions that help the LLM answer the same questions better. Suggestions should provide specific and detailed guidance on logical thinking steps, required knowledge, and abilities, ensuring the LLM can answer correctly for the same questions. Output in the following JSON format: {{   ”flaws_knowledge”: ”The lack of background knowledge.”,   ”flaws_capability”: ”The flaws in logic and capability.”,   ”comprehensive_performance”: ”The Comprehensive performance of all questions.”,   ”suggestions”: ”Suggestions that help the LLM answer questions bette” }}
@@
Prompt for Querying LLM with suggestions Please complete the following question: [question]: { question } In your previous responses to these questions, the interviewer has provided the following suggestions for you to help you answer better: [suggestions]: { suggestions } Please consider the above [suggestions], and answer the above [question], in the following JSON format: {{”answer”: ”Your answer”}}

### v4

@@ Appendix B Analysis and Proofs of Adaptive Difficulty-Control Mechanisms
$\textbf{C1.}\quad avg_{s}^{\prime}=\frac{avg_{s}N+c_{it}}{N+1}>avg_{s},$ (1)
$\quad\text{holds true always when}\ \ avg_{s}\leq t_{(i,i+1)}$
$\textbf{C2.}\quad\exists\ \ avg_{s0},\,N_{0},$
$\quad\text{s.t.}\ avg_{s}^{\prime}=\frac{avg_{s0}N_{0}+c_{it}}{N_{0}+1}>t_{(i,i+1)}$
@@ Appendix C Construction of Context Graph
$N=(e\,,\,\mathcal{C}\,,\,\mathcal{D})\ ,$ (5)
$\quad\text{where}\ \mathcal{C}=\{c_{1},c_{2}.\}\ ,\ \mathcal{D}=\{d_{1},d_{2},$.\}
@@ D.1 Details of Datasets
MultiHop-RAG(Tang and Yang, 2024) consists of phrase Q&A queries, their ground truth answers, and the associated supported evidence constructed from news articles published between September and December 2023. The background knowledge texts of MultiHop-RAG are provided as supporting evidence along with the questions.
D.2 Preprocessing of Datasets
To verify JudgeAgent’s ability to evaluate knowledge deficiencies in target models, we remove the background knowledge of the questions from MedQA and MultiHopRAG during evaluation. Otherwise, it would only test the target’s reading comprehension ability rather than knowledge deficiencies. In contrast, we provided the background text when using QuALITY for evaluation, as the questions in QuALITY are highly dependent on the text content, and many of the texts are fictional narratives. Therefore, we use QuALITY to validate the JudgeAgent’s effectiveness in guiding comprehension and reasoning rather than discovering knowledge deficiencies.
@@ Appendix E Additional Experiment Analysis
E.1 Efficiency of JudgeAgent
Target MedQA MultiHopRAG
avg.token avg.time avg.token avg.time
GLM4-Flash 1558 7.24 2424 23.38
Qwen3 1909 9.95 3453 30.96
Table 5: Efficiency of JudgeAgent. avg.tokens and avg.time respectively denote the average tokens and time (seconds) cost of JudgeAgent per seed question.
We calculated the time and token costs incurred by JudgeAgent during multi-turn interactions with GLM4-Flash and Qwen3 as target models, calculated as averages based on the number of questions in the benchmark selected during the Benchmark Grading stage. The results are shown in Table 5.
E.2 Stability of JudgeAgent
Target MedQA
[…]
Target MultiHopRAG
[…]
Table 6: The experiment results with standard deviation calculated based on 5 different random seeds.
To investigate the stability of JudgeAgent’s results, this section presents additional experiments conducted under 5 different random seeds, with the standard deviation calculated. Experiments are conducted using Qwen3 and GLM4-Flash as target models on the MedQA and MultiHop-RAG benchmarks. The results are shown in Table 6.
The experiment results demonstrate that the results across multiple random seeds remain stable and are consistent with the ranking of results presented in the main results in Table 2. This stability benefits from JudgeAgent’s context-graph control and knowledge-driven synthesis pathway for evaluation questions. These mechanisms endow the generated questions and the evaluation feedback with controllability and stability.
E.3 The Resilience Against Data Contamination
Can JudgeAgent mitigate the challenges of data contamination in static benchmarking paradigms? In this section, we simulate a scenario where the static benchmarking evaluation paradigm suffers from data contamination by deliberately exposing the evaluation questions to LLMs during their training process. We selected Llama3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2.5-7B-Instruct as base models, and constructed supervised fine-tuning (SFT) data from the MedQA and MultiHop-RAG benchmarks, which are intended for evaluation in the main experiments. These training data were used to fine-tune the selected base models. By comparing the performance differences between the original base models and the fine-tuned models on both the static benchmark questions and the extended questions generated by JudgeAgent, we analyze and verify the resilience of JudgeAgent and its derivative JudgeAgent against data contamination.
@@ E.3 The Resilience Against Data Contamination
Table 7: The results for validating the resilience against data contamination. base means the performance on base static benchmark questions, and @K means extended questions at the K-th iteration.
@@
In contrast, when evaluated on extended questions generated by JudgeAgent, the fine-tuned models and base models show little difference in performance. Additionally, fine-tuning even resulted in a decline on the MedQA benchmark. These results suggest that under the JudgeAgent dynamic evaluation paradigm, the generated questions maintain the validity of the evaluation, even when the original questions have been exposed to the target model. Furthermore, under the same setting of LLM and benchmark, there is a marked gap between the performance gain on base questions and extended questions after fine-tuning, demonstrating that JudgeAgent exhibits considerable resilience to data contamination.
Target ACC1 ACC2d ACC2f
Qwen3 91.71 99.18 96.38
GLM4-Flash 80.09 98.37 92.82
GPT-4.1 84.97 98.28 92.44
Table 8: The comparative results on MedQA of directly providing correct answers (ACC2d) and providing feedback from JudgeAgent (ACC2f). ACC1 represents the accuracy of target models before evaluation. All values are percentages.
Additionally, we conducted a comparative experiment providing the target models with both standard evaluation feedback from JudgeAgent and evaluation feedback containing the correct answers as cheating information, investigating the risk of data contamination during the evaluation. The results are presented in Table 8. As can be observed, when the correct answers are directly provided, nearly all LLMs achieve almost perfect accuracy. This further illustrates the risk of data contamination in existing static benchmarks. It also demonstrates that the evaluation feedback from JudgeAgent does not contain cheating information tailored specifically to the seed questions
E.4 Quality of Evaluation Questions
We use Dingo [^3] https://github.com/MigoXLab/dingo to check the factuality of evaluation questions generated by JudgeAgent in multi-turn interactions. The results are shown in Table 9, $N_{eq}$ represents the number of generated questions in Staget Interactive Extension.
Datasets $N_{eq}$ avg. fact ratio fact check passed
$N_{passed}$ passed ratio
MedQA 3819 97.93% 3680 96.36%
MultiHop-RAG 7668 94.38% 7185 93.70%
Table 9: The fact check results by Dingo.
As shown in Table 9, in Dingo’s fact-checking task, the extended questions from JudgeAgent maintain a high fact ratio and detection pass ratio (using Dingo’s default fact ratio threshold of 80%). The average fact ratio is above 94%, while the pass ratio exceeds 93% for both MedQA and MultiHop-RAG. The results demonstrate that the most generated questions exhibit strong factuality, ensuring that the dynamic evaluation results are based on high-quality questions, and the probability of an evaluated LLM being incorrectly judged as wrong when its answer is correct remains low.
As for the diversity of evaluation questions, the experiment results of data contamination in Table 7 in Appendix E.3 can indirectly demonstrate that contaminated models perform poorly on synthesized questions, which indirectly confirms that synthesized questions differ from seed questions and that their diversity is assured.
E.5 Cross Validation of the Evaluation Suggestions
Do the suggestions provided by JudgeAgent only take effect for the seed questions? To address this question, we designed cross-validation experiments from two perspectives.
First, considering that the evaluation suggestions are derived from a comprehensive evaluation of the responses to both the seed questions and their extended questions, we evaluated and compared the accuracy improvement of the suggestions on the extended questions versus the seed questions, aiming at verifying that the suggestions do not contain "cheating information" specific to seed questions in the scenario after evaluation. The results are shown in Figure 6.
@@ E.5 Cross Validation of the Evaluation Suggestions
Secondly, we transferred the evaluation suggestions derived from seed questions to other questions with related knowledge concepts in the benchmark, to verify the effectiveness of the suggestions within the same knowledge domain. Specifically, we categorized questions based on the knowledge entities they contain, formalized as follows: given the context graph $\mathcal{G}$, for two questions $q_{1}$ and $q_{2}$ with knowledge entities $E_{1}=\{e|e\in q_{1}\land e\in\mathcal{G}\}$ and $E_{1}=\{e|e\in q_{2}\land e\in\mathcal{G}\}$, if $E_{1}$ and $E_{2}$ have a non-empty intersection, then $q_{1}$ and $q_{2}$ are considered related questions. In this experiment, the suggestions for a question were constructed from the suggestions of its related questions, excluding suggestions from the question itself, to assess and validate the transferability of the suggestions provided by the JudgeAgent. We screened out questions without relevant questions and those with only relevant questions based on non-knowledge entities, such as male, female, 2 years, etc. The results are shown in Table 10.
@@
First, we analyze the difference in the effectiveness of evaluation suggestions on seed questions versus extended questions. As shown in Figure 6, suggestions effectively improve performance for both seed questions and extended questions, with a relatively small gap in the degree of improvement. Notably, when Qwen3 and GPT-4.1 are used as target models, the first round of Qwen3 and the second round of GPT-4.1 exhibit even greater improvement than seed questions. These results indicate that although the suggestions only supplement knowledge for several key concepts, such as the case in Figure 8, such concise suggestions can still benefit both seed and extended questions, demonstrating that JudgeAgent is capable of identifying and addressing knowledge gaps in the target model, rather than simply providing "cheating information" specific to seed questions.
@@
Next, we analyze the difference in the effectiveness of suggestions on questions that share the same knowledge concepts. As shown in Table 10, compared to their effectiveness on seed questions, the suggestions exhibit a slight decrease when applied to related questions, as indicated by a decline in CR and $\Delta$, and an increase in CtE. But the difference is minor, and the improvement remains notable, suggesting that the suggestions can be effectively transferred to other questions involving the same knowledge concepts. This further demonstrates that JudgeAgent can provide suggestions that do not simply serve as “cheating information” specific to seed questions.
However, the slight decline in evaluation effectiveness also indicates that the knowledge guidance provided in the suggestions is not fully aligned with the related questions. This may be because the overlapping entities between these related questions and seed questions do not correspond to core knowledge concepts. For example, suggestions centered on "blood type" may be transferred to a question where "serum" is the core knowledge concept, resulting in a partial mismatch.
The above experiment results demonstrate that the evaluation suggestions provided by JudgeAgent are not only applicable to seed questions but can also be transferred to other questions that share relevant core knowledge concepts.
E.6 Supplementary Parameter Analysis
What is the impact of batch size in the Benchmark Grading stage on the evaluations of the JudgeAgent? In Benchmark Grading stage, questions are divided into batches to comprehensively assess the target’s capabilities in a base level. These batches are also the basic units for question extension and evaluation feedback. Given a fixed rounds, the batch size is inversely proportional to the number of batches, extended questions, and evaluation suggestions, thereby influencing the time and resource consumption of the entire evaluation process. Can the batch size be maximized to reduce resource consumption while maintaining the effectiveness of evaluations? To address this question, we conducted a parameter analysis experiment, examining the evaluation effectiveness and time consumption under different batch sizes. The results are shown in Figure 7.
@@ Appendix G Prompts
We show the detailed prompt for Entity Extraction in Prompt 11, the prompt for generating extended questions in Prompt 12, the prompt for evaluating the performance of the target LLM in Prompt 13, and the prompt for querying the target LLM with suggestions in Prompt 14.
Prompt for Entity Extraction Please identify and label the entities in the following multiple sentences, and return the entity labeling results for each sentence. The results for each sentence should be independent, in JSON format, containing the sentence number, sentence text, and the list of recognized entities (including entity text, type, and position). Return format is a dictionary, with only one key ’labeled_data’, and the value is a list, each element is a dictionary containing the sentence text and the entity list. {{   "labeled_data":   [     {{"text":"Sentence 1", "entity_list": [{{"entity_text": "", "entity_type": ""}}]}},     {{"text":"Sentence 2", "entity_list": [{{"entity_text": "", "entity_type": ""}}]}},     …   ] }} Notice that "text" should be only the sentence, not the whole article. Sentence list: { Sentences }
@@
Prompt for Generating Extended Questions As an interviewer, you are tasked with designing questions based on the provided texts. Your role involves crafting questions and correct answers that fulfill the following criteria: 1. **Focus on the Entity**: Ensure all questions consistently center around the specified entity from the article. 2. **Ensure Accuracy and Conciseness of Answers**: Verify that the provided answer is both correct for your designed question within the context and logic of the given text fragments, and ensure the answer is sufficiently concise—presented as a word or phrase, avoiding redundancy. 3. **Conform to difficulty requirements**: You need to design questions for the required difficulty levels, with specific requirements as follows:   (1). [easy]: **Encourage Knowledge Memorization**, design questions that assess whether respondents have memorized relevant knowledge. Create questions by directly extracting and blanking out content from the given passage.   (2). [medium]: **Encourage Knowledge Comprehension**: Design questions that prompt respondents to dissect and comprehend concepts involved in the topic. Avoid assessing only superficial knowledge retention.   (3). [hard]: **Encourage Knowledge Deep Analysis**: Design questions that prompt respondents to engage in deep thinking and analysis. Avoid merely testing knowledge recall or conceptual comprehension; do not simply extract fragments from the given passage to create fill-in-the-blank items. Encourage respondents to focus on entities within the question and employ logical skills for complex reasoning. Here are examples: { Examples } Now, given the following text fragments: { context } Based on the provided texts, please generate questions by following the requirements above and referencing the examples. Output in the specified JSON format below: {{   "generated_question":   [     {{       "question": "Generated Question",       "answer": "Correct Answer of Generated Question"     }},   …   ] }}
@@
Prompt for Evaluation The following is the performance of an LLM in answering a series of questions: { list of questions, correct answers, and LLM’s answers } Please evaluate and analyze the interviewee’s performance based on the above performance using concise language from the following perspectives, and provide suggestions that help the LLM answer the same questions better. Suggestions should provide specific and detailed guidance on logical thinking steps, required knowledge, and abilities, ensuring the LLM can answer correctly for the same questions. Output in the following JSON format: {{   "flaws_knowledge": "The lack of background knowledge.",   "flaws_capability": "The flaws in logic and capability.",   "comprehensive_performance": "The Comprehensive performance of all questions.",   "suggestions": "Suggestions that help the LLM answer questions bette" }}
@@
Prompt for Querying LLM with suggestions Please complete the following question: [question]: { question } In your previous responses to these questions, the interviewer has provided the following suggestions for you to help you answer better: [suggestions]: { suggestions } Please consider the above [suggestions], and answer the above [question], in the following JSON format: {{"answer": "Your answer"}}